# Cocoa: Co-Planning and Co-Execution with AI Agents

### v3

@@ 4.3. Interleaving Co-Planning and Co-Execution
In addition to supporting co-planning and co-execution independently, Cocoa is designed to support synergistic interleaving of the two. As Nóírín co-executes the plan with the agent, she encounters a relevant author that the agent identified and wishes to further explore the author’s papers. She edits the following step’s plan description to satisfy this. She reruns that step and receives a new batch of papers, and the agent automatically reruns the rest of the plan with the updated output. She is also not satisfied with the agent’s interpretation of common themes and believes that she can surface deeper insights by reading the papers herself, so she toggles that step to be a user step. After reading the papers in detail, she returns to her document and jots down her insights. She now finds the next step repetitive, so she deletes it. Overall, her use of Cocoa is highly iterative, and she smoothly transitions between co-planning and co-execution.
Figure 10. Users can create plans at multiple locations in their document and have the agent tackle them in parallel. Plans can be collapsed to only reveal essential information about their operating status.
[…]
An illustration of three plans in a document, with two of them collapsed so that only essential information about their operating status is shown without details about the individual steps.
So far, Nóírín has just been interacting with one plan. While waiting for the agent to complete a series of steps, she can collapse the plan to hide the steps and only reveal essential information about the plan’s status. She identifies a couple more areas of her document that merit further exploration and invokes the agent on them. Soon, she is co-planning and co-executing with multiple agents on different parts of her document in parallel. This is not cognitively burdensome for her as she only has to attend to one plan at a time, since the others are all collapsed and running in the background (Figure 10). Finally, the document in Cocoa supports real-time multi-user editing, so Nóírín’s collaborators can also leverage the agent and interact with plans.
@@ 4.4. Implementation Details
4.4.1. Technical stack
Cocoa is implemented as a web application with a Next.js and TypeScript frontend communicating with a Flask backend. The frontend uses the Tiptap [^3] https://tiptap.dev/. framework for the main document editor. Each document supports synchronous collaboration via Hocuspocus [^4] https://tiptap.dev/docs/hocuspocus/. and all changes are auto-saved to Tiptap Cloud, which also serves as storage for all participant documents. Interactive components within the document editor are implemented as custom Tiptap extensions in TypeScript.
4.4.2. Underlying LLM Agent
The Flask backend orchestrates LLM activity with calls to GPT-4o (scaffolded by LangChain [^5] https://www.langchain.com/). To optimize for speed, simple actions such as plan generation or text summarization involve a direct call to GPT-4o. For plan steps that required access to the scientific literature, we created a custom tool-calling LLM agent powered by GPT-4o that had two tools that allowed it to access the public Semantic Scholar API [^6] https://www.semanticscholar.org/product/api, which covers paper/author/topic search, and the Ask this Paper [^7] https://www.semanticscholar.org/product feature for paper summarization and question-answering. While GPT-4o provides APIs for managing multi-turn conversation context, the system manages its own context, which includes previous plan steps and their outputs, and calls GPT-4o’s single-turn completion API. All prompts used are provided in Appendix I.
4.4.3. Guidance for plan generation
We use findings from our formative study to guide the agent’s generation of initial plans. In our formative study, we had asked participants to write down brief plans when interacting with our probe (see Appendix C). Based on participants’ preferences of which steps they prefer assigning to an AI vs. keeping for themselves (F3 from our formative study), we organized these steps into agent and user steps. We then wrote examples of how these steps are composed into plans for tackling particular questions, once again drawing from plans participants wrote in the probe activity. These examples were provided for in-context learning in Cocoa’s system prompt for plan generation (see prompts in Appendix I).
We also enable Cocoa to learn from interactive plans the user has previously created and edited. When the agent is invoked, Cocoa collects the existing plans in the document and adds them to the in-context learning examples on-the-fly. This allows the user’s co-planning efforts to be reused for similar requests elsewhere in the document.
@@ 5. Lab Study
We conducted a within-subjects task-based lab study with 16 researchers to evaluate Cocoa and its interactive plan design against a strong chat baseline. Specifically, our study aimed to address the following research questions:
@@
    How does Cocoa compare to our chat baseline for ease of use, steerability, and general utility in research project documents?
@@
    When did researchers prefer interacting with an AI agent through interactive planning versus our chat baseline?
@@ 5.2. Baseline System
The baseline system (Figure 11) was a chat interface powered by the same LLM agent used in Cocoa: it used GPT-4o with access to the same tools for completing literature-related tasks using a combination of Semantic Scholar and standard LLM capabilities (summarization, accessing knowledge stored in weights, etc.). While we could integrate interactive plans within chat and vice versa (more on this in Section 8.1), we opted not to for a cleaner comparison of the two design patterns. The design of the chat interface closely resembled that of popular LLM chatbots (e.g., ChatGPT, Claude). During the study, participants positioned the chat interface beside the baseline document editor, which is the same editor that Cocoa uses, except we removed the option to invoke the agent to eliminate all co-planning and co-execution interactions.
@@ 5.3. Research Task and Study Procedure
The first author (study facilitator) conducted 1:1 studies with the 16 participants over Google Meet between October and December 2024. Each study was 90 minutes in length. The study started off with introductions and a brief overview, followed by the two tasks, which were counterbalanced across participants. Detailed procedures for each condition can be found in Appendix [x]. After each condition, participants filled out a short evaluation form with 5-point Likert scale questions for that condition (see Appendix H). The task-based portion of the study took up 75 minutes. The study concluded with a semi-structured interview that lasted around 15 minutes. Participants were asked to reflect on their experiences across the two systems and discuss the pros and cons of both. They were also asked about particular decisions the study facilitator observed when using each system.
Participants received a \$75 USD honorarium upon completion of the study. All studies were recorded and transcribed by Google Meet. This study was reviewed and approved by our organization’s internal IRB.
@@ 5.4. Data Analysis
The first author also coded all video recordings from the study to note the timestamps at which participants engaged and disengaged in a particular interaction. In Cocoa, these interactions were co-planning (see 4.1), co-execution (see 4.2), and output inspection (when the participant passively inspected the system’s outputs). The closest equivalent interactions were used as codes in the baseline: prompting for planning (instructing the agent to perform an action without any attempts to modify the agent’s earlier outputs), prompting for editing [^9] Since there were no user steps in the baseline, the closest equivalent interaction is attempting to edit the agent’s output via prompting. (instructing the agent to modify or only consider a subset of its output in future actions), and output inspection. During video coding, the first author also captured the number of input items (items generated as part of the output of a plan step or conversation turn) and output items (items fed as context into the next plan step or conversation turn) for all plan steps or turns that output discrete items (i.e., papers, authors, topics, text entities).
@@ 6. Lab Study Results
A high-level overview of participants’ interactions with Cocoa and the baseline is depicted in Figure 12. In both conditions, output inspection occupied the most time out of the coded interactions, but was lengthier in the baseline than Cocoa (54.1% vs. 33.6% of the approximately 25-minute session). A paired t-test found this difference to be significant ($t(15)=3.95,p<0.001$). Participants also engaged in co-execution in Cocoa significantly more often than the closest equivalent interaction—output editing via prompting—in the baseline (15.2% vs. 3.1%, $t(15)=4.29,p<0.001$). They spent slightly more time on co-planning (21.5% vs. 17.4%), although this was not significant ($t(15)=1.01,p=0.33$). Our analysis indicates that, when provided with more affordances to interactively modify the agent’s outputs, users will take advantage of those opportunities for active engagement and shift away from passive output consumption.
@@
Could more output editing activity mean that participants are less satisfied with Cocoa’s outputs? To answer this, we analyzed the output-to-input item ratio for each round of interaction—where a round is a plan step in Cocoa or conversation turn in the baseline—across all participants. The output-to-input ratio tells us what percentage of items generated by a round are carried over as context to be used in the next round. As can be seen in Figure 13, the ratio is much higher in Cocoa ($M=97.5,\sigma=23.0$) compared to the baseline ($M=0,\sigma=33.0$). This difference is significant per a Mann-Whitney U Test ($U=1722.0,p<0.001$). These results suggest that the flow of data within the two systems differ fundamentally: the flow of items from one plan step to the next is rather fluid in Cocoa where much results were carried over to the next steps in the plan, but disjoint in the baseline where most or all of the results were unused. This suggests that participants, when using the baseline system, more often focused on few items in the output, or pivoted to a different task entirely. Thus, there is no indication that increased output editing activity is due to dissatisfaction with system outputs.
Figure 13. The output-to-input ratio in Cocoa and the baseline, across all rounds of interaction from all participants ($n=16$). *** indicates $p<0.001$ via a Mann-Whitney U Test.
[…]
A boxplot depicting the output-to-input ratio in Cocoa and the baseline. Cocoa is shown above the baseline on the y-axis and has a median of 97.5, whereas the baseline has a median of 0. The output-to-input item ratio is shown on the x-axis from 0 to 100. The difference between the two conditions is highly significant.
In the rest of this section, we present quantitative and qualitative analyses of participants’ data from our user study. We group these results by our research questions from the start of Section 5.
@@ 6.1. Steerability, Ease of Use, and Utility (L1)
We denote the median rating of our baseline and Cocoa as $M_{b}$ and $M_{c}$, respectively, and the Wilcoxon test statistic as $W$.
Introducing novel interactive systems with additional affordances can often lead to higher effort when using the systems as a trade-off for better utility (Shneiderman and Maes, 1997; Hutchins et al., 1985). However, based on participants’ post-task ratings on a 5-point Likert scale, there was no significant difference in perceived ease of use between the baseline and Cocoa with 16 participants ($M_{b}=4.5,M_{c}=4,p=1.000,W=7.50$). At the same time, we observed a significant difference that Cocoa provided better steerability. In response to the question “I could easily steer the system towards doing something helpful,” participants rated Cocoa higher than the baseline to a significant degree ($M_{b}=3,M_{c}=4,p=0.005,W=0$), [^10] Here, we set the significance threshold to $\alpha=0.05/3=0.015$ using Bonferroni correction. Our result is significant post-correction: $p=0.005<0.015$. with $W=0$ indicating that all participants rated Cocoa’s steerability as greater than or equal to that of our baseline. In sum, these results suggest that by taking inspiration from familiar interaction patterns of computation notebooks and text editors, Cocoa provided richer affordances, and does so without sacrificing ease of use to better support human-AI collaboration compared to the simple chat interface used in the baseline condition (Figure 14).
@@
Additionally, the nature of the task may have influenced perceived utility, as Cocoa might be better suited for certain tasks, while chat may work better for others. Qualitative insights from post-task interviews and participants’ think-aloud sessions provide insights into the task suitability of interactive plans (see Section 6.2.1).
6.1.1. Co-planning afforded steerability
@@
Qualitative results based on think alouds and post-interviews showed that participants found various co-planning features provided by Cocoa helped improve the steerability of the agent. Many thought that the ability to compose, edit, and rerun plan steps allowed them to better “fine-tune” the agent’s outputs and their own research process. P1 was inspired to “use [plan steps] as building blocks” to create a custom workflow. P7 and P15 both appreciated the ability to edit and rerun steps, which enabled them to quickly “organize thoughts and iterate on ideas” (P7); P15, specifically, iterated on “paper search” step, reran it to cover papers from venues that they preferred, and was able to obtain more relevant final outputs with running the rest of the steps in the plan. Another strategy was “backtracking” to a specific step and editing it which P6 found to be much more helpful than digging through a chat conversation: “With chat, when it does something wrong there’s not really an easy way to fix it because there’s no concrete steps that it’s following, whereas [with Cocoa], I can go back and be like, let me have it do something else here.”
@@
Besides iterating and steps by editing and rerunning them, we also observed participants adding and removing agent-suggested steps to improve the plans. For example, P12 added a step at the end of their plan to have the agent suggest some potential next steps to pursue after seeing its paper summaries, while P5 removed a less relevant intermediate step on searching for papers related to autonomous vehicles to avoid distractions from this tangential topic.
6.1.2. Co-execution afforded steerability
Participants appreciated the flexibility to choose between stepwise and continuous plan execution. Many chose stepwise execution because it allowed them to better control the overall process and prevent error propagation. P4 and P11 both wanted to verify and manually curate the list of papers returned by the agent from a paper search step before proceeding to subsequent steps. P11 specifically cites resource efficiency as a practical reason: “If I update [outputs of] the first or second step it would need to rerun the following steps. I just want to reduce some API usage.” P10 similarly mentioned that they would “waste energy” running subsequent steps if they detected errors in initial steps. Others preferred continuous execution. P8 and P9 thought continuous execution was more efficient because it parallelized human and AI efforts: “Running all is a little bit of parallelization in my head where I can get the next thing rolling, and if I end up not having to edit it, that means we’re already moving ahead” (P8). P15 stated that while they opted for continuous execution, their choice depended on the familiarity with the task: “if it were a topic that I felt like I didn’t already know what type of literature I would be looking for or what the output should look like, I would go through them individually and maybe use it as more of a literature exploration tool. But in this case, I already know in general what types of papers I want to see it come up with.” This underscores the importance of providing users with a choice of execution method to cater to researchers’ contextual needs and preferences.

### v4

@@ 4.3. Interleaving Co-Planning and Co-Execution
In addition to supporting co-planning and co-execution independently, Cocoa is designed to support synergistic interleaving of the two. As Nora co-executes the plan with the agent, she encounters a relevant author that the agent identified and wishes to further explore the author’s papers. She edits the following step’s plan description to satisfy this. She reruns that step and receives a new batch of papers, and the agent automatically reruns the rest of the plan with the updated output. She is also not satisfied with the agent’s interpretation of common themes and believes that she can surface deeper insights by reading the papers herself, so she toggles that step to be a user step. After reading the papers in detail, she returns to her document and jots down her insights. She now finds the next step repetitive, so she deletes it. Overall, her use of Cocoa is highly iterative, and she smoothly transitions between co-planning and co-execution.
When designing Cocoa, we were inspired by the computational notebook paradigm supported scientists via tight coupling of code and output, enabling rapid iteration and steering of complex scripts. To summarize this connection, we map Cocoa’s features to those in computational notebooks in Table 1.
Cocoa Feature Computational Notebook Feature
Editable plan steps that can be composed into a larger, more complex task. Code cells that can be composed into a larger, more complex program.
Users can choose between stepwise and continuous execution. Users can choose between running individual cells or run all.
Users can view and improve step outputs within the plan. Users can view code outputs within the notebook.
Users can flexibly switch between improving the plan and executing the plan steps. Users can flexibly switch between improving the code and executing code cells.
Plan steps assigned as user steps pause execution so users can perform tasks themselves. Break points pause execution so users can examine current code outputs; all cells are executed programmatically.
Completed plan can serve as documentation of agent & human work. Completed notebook can serve as documentation of analyst’s work.
Table 1. Main connections between features in Cocoa and those in computational notebooks.
A 2-column table with headings Cocoa Feature and Computational Notebook Feature. There are 3 rows, where each row contains a description of the feature in Cocoa and computational notebooks.
@@ 4.4. Implementation Details
Cocoa is implemented as a web application with a Next.js and TypeScript frontend communicating with a Flask backend. The frontend uses the Tiptap [^3] https://tiptap.dev/ framework for the main document editor. Each document supports synchronous collaboration via Hocuspocus [^4] https://tiptap.dev/docs/hocuspocus/ and all changes are auto-saved to Tiptap Cloud, which also serves as storage for all participant documents. Interactive components within the document editor are implemented as custom Tiptap extensions in TypeScript.
The Flask backend orchestrates LLM activity with calls to a custom tool-calling LLM agent [^5] https://platform.openai.com/docs/guides/function-calling powered by GPT-4o and scaffolded by LangChain. [^6] https://www.langchain.com/ Our agent is designed to assist researchers through literature-grounded response generation. To do this, we implemented tools that give our agent access to millions of research papers via publically available Semantic Scholar APIs. [^7] https://www.semanticscholar.org/product/api This API is powered by The Semantic Scholar Open Data Platform (Kinney et al., 2023) which consists of a state-of-the-art PDF extraction and knowledge graph normalization pipeline. This gave our agent access to both the full text of millions of research papers and also their rich metadata, such as authors, fields of studies, and publication venues (Kinney et al., 2023). More implementation details can be found in Appendix D.
@@ 5. Lab Study
We conducted a within-subjects task-based lab study with 16 researchers to evaluate Cocoa against a custom chat baseline. Specifically, our study aimed to address the following research questions:
@@
    How does Cocoa compare to our chat baseline for ease of use, steerability, and general utility [^8] General utility is a composite metric created from Q1, Q2, and Q4 from Appendix H. in research project documents?
@@
    How does interleaved co-planning and co-execution contribute to ease of use, steerability, and general utility in Cocoa, if there is a significant improvement over the baseline?
@@ 5.2. Baseline System
As our primary research question is to better understand the costs and benefits of Cocoa’s co-planning and co-execution affordances, the baseline system (Figure 7) was a common chat interface powered by the same underlying LLM agent: a custom tool-calling agent powered by GPT-4o with access to the same state-of-the-art tools for searching, analyzing, summarizing, and comparing literature across millions of papers on Semantic Scholar (Kinney et al., 2023), in addition to standard LLM capabilities. See D for more details about the agent. The design of the chat interface closely resembled that of existing AI-powered research systems (e.g., Assist mode in Scite (Scite, 2025), Google Scholar Labs, various deep research systems), and popular LLM chatbots (e.g., ChatGPT, Claude). During the study, participants positioned the chat interface beside the baseline document editor, which is the same editor that Cocoa uses, except we removed the option to invoke the agent to eliminate all co-planning and co-execution interactions.
We chose this baseline because it mimicked the dominant interaction paradigm implemented in AI-powered research tools at the time of our study and allowed for strong comparisons on aspects like ease of use. We recognize that variations of this baseline, such as forcing the agent to show an editable plan before execution (e.g., (He et al., 2025)), are also strong baselines that can help further isolate the factors contributing to our results. We discuss ablations Cocoa as alternative baselines in Section 9.
@@ 5.3. Research Task and Study Procedure
The first author (study facilitator) conducted 1:1 studies with the 16 participants over Google Meet between October and December 2024. Each study was 90 minutes in length. The study started off with introductions and a brief overview, followed by the two tasks, which were counterbalanced across participants. Detailed procedures for each condition can be found in Appendix G. After each condition, participants filled out a short evaluation form with 5-point Likert scale questions for that condition (see Appendix H). These questions were inspired by the System Usability Scale (Brooke and others, 1996) but adapted for the research context.
The task-based portion of the study was 75 minutes. The study concluded with a semi-structured interview that lasted around 15 minutes. Participants were asked to reflect on their experiences across the two systems and discuss the pros and cons of both. They were also asked about particular decisions the study facilitator observed when using each system.
Participants received a \$75 USD honorarium upon completion of the study. All studies were recorded and transcribed by Google Meet. This study was reviewed and approved by the Allen Institute of AI’s IRB.
@@ 5.4. Data Analysis
The first author also coded all video recordings from the study to note the timestamps at which participants engaged and disengaged in a particular interaction. In Cocoa, these interactions were co-planning (see 4.1), co-execution (see 4.2), and output inspection (when the participant passively inspected the system’s outputs). The closest equivalent interactions were used as codes in the baseline: prompting for planning (instructing the agent to perform an action without any attempts to modify the agent’s earlier outputs), prompting for editing [^10] Since there were no user steps in the baseline, the closest equivalent interaction is attempting to edit the agent’s output via prompting. (instructing the agent to modify or only consider a subset of its output in future actions), and output inspection.
We established these equivalences based on the functional purpose of the interactions: co-planning and prompting for planning both instruct the agent to produce new outputs, while co-execution and prompting for editing aim to improve an existing output. We note that there is not perfect correspondence between interactions across the two systems. For example, co-execution is not only about refining agent outputs, but also giving users agency to complete tasks they assigned for themselves. We acknowledge this by shading the interactions in Cocoa and the baseline with slightly different colours in Fig. 8. However, because co-planning and co-execution subsume the functional roles of their baseline counterparts in the context of our study, we consider them appropriate to compare.
@@ 6. Lab Study Results
A high-level overview of participants’ interactions with Cocoa and the baseline is depicted in Figure 8. In both conditions, output inspection occupied the most time out of the coded interactions, but was lengthier in the baseline than Cocoa (54.1% vs. 33.6% of the approximately 25-minute session). A paired t-test found this difference to be significant ($t(15)=3.95,p<0.001$). Participants also engaged in co-execution in Cocoa significantly more often than the closest equivalent interaction—output editing via prompting—in the baseline (15.2% vs. 3.1%, $t(15)=4.29,p<0.001$). They spent slightly more time on co-planning (21.5% vs. 17.4%), although this was not significant ($t(15)=1.01,p=0.33$). Our analysis indicates that, when provided with more affordances to interactively modify the agent’s outputs, users will take advantage of those opportunities for active engagement and shift away from passive output consumption. That is, users could interleave co-planning and co-execution to steer the agent through tighter feedback loops, rather than relying on slower cycles of trial-and-error prompting and output inspection.
We also computed performance metrics for both co-planning and co-execution from coded interactions and system logs. These included the number of step edits, reassignments, and output edits across all participants. Table 2 displays our results. Overall, 32.6% of steps were edited and 17.4% were reassigned. This shows that most plan steps proposed by the agent (the content of the step as well as the assignment) were accepted as-is. During execution, the agent surfaced an average of 12.9 discrete output items (i.e., papers, authors, topics, text entities) per step to users. On average, 2.1 (16.3%) were deleted by participants, suggesting that the system had reasonable precision.
However, these metrics alone may not paint a holistic picture about Cocoa’s use. To more deeply understand the context and rationale behind users’ actions, we analyzed videos from our user studies (Fig. 8) alongside qualitative transcripts from participants thinking aloud (Section 6.2).
Metric Count (%, if
applicable)
Total steps completed across all plans 86
Step edits by user 28 (36.5%)
Step reassignments by user 15 (17.4%)
Step additions by user 12 (14.0%)
Step deletions by user 4 (4.7%)
Replanning triggered by system 14
Avg. output items (i.e., papers, authors, topics, text entities) per step 12.9
Avg. output items deleted per step 2.1 (16.3%)
Table 2. Cocoa usage metrics from our lab study.
A 2-column table with headings Metric and Count. There are 8 rows, where each row contains a metric with its corresponding count and percentage, if applicable.
@@
In the rest of this section, we present quantitative and qualitative analyses of participants’ data from our user study. We group these results by our research questions from the start of Section 5. We denote the median rating of our baseline and Cocoa as $M_{b}$ and $M_{c}$, respectively, and the Wilcoxon test statistic as $W$.
@@ 6.1. Steerability, Ease of Use, and Utility (L1)
Introducing novel interactive systems with additional affordances can often lead to higher effort when using the systems as a trade-off for better utility (Shneiderman and Maes, 1997; Hutchins et al., 1985). However, based on participants’ post-task ratings on a 5-point Likert scale, there was no significant difference in perceived ease of use between the baseline and Cocoa with 16 participants ($M_{b}=4.5,M_{c}=4,p=1.000,W=7.50$). At the same time, we observed a significant difference that Cocoa provided better steerability. In response to the question “I could easily steer the system towards doing something helpful,” participants rated Cocoa higher than the baseline to a significant degree ($M_{b}=3,M_{c}=4,p=0.005,W=0$), [^11] Here, we set the significance threshold to $\alpha=0.05/3=0.015$ using Bonferroni correction. Our result is significant post-correction: $p=0.005<0.015$. with $W=0$ indicating that all participants rated Cocoa’s steerability as greater than or equal to that of our baseline. In sum, these results suggest that Cocoa affords interleaving co-planning and co-execution via rich interactive affordances without sacrificing ease of use, when compared to the simple chat interface used in the baseline condition (Figure 9).
@@
6.2. Interleaving Co-Planning and Co-Execution Created an Effective Steering Loop (L2)
6.2.1. Outputs from co-execution drove re-planning
@@
Qualitative results based on think alouds and post-interviews suggest that Cocoa was valuable not only because it enabled human-agent co-planning or co-execution, but because interleaving the two meant that co-execution outputs could be used to inform collaborative re-planning. Rather than treating the plan as a static artifact, participants used co-execution to find interesting and potentially productive directions to steer the agent towards. For example, P7 and P15 both used outputs from agent steps as a signal to “organize thoughts and iterate on ideas” (P7). P15 specifically iterated on an intermediate “paper search” step by re-executing it to cover papers from venues that they preferred. This led to subsequent plan updates that ultimately resulted in a more relevant final output. Similarly, P6 observed that “backtracking” to a specific step and editing it was more helpful than digging through a chat conversation: “With chat, when it does something wrong there’s not really an easy way to fix it because there’s no concrete steps that it’s following, whereas [with Cocoa], I can go back and be like, let me have it do something else here.”
@@
Besides modifying intermediate steps, participants also added steps at the end of the plan to assign follow-up investigations, or deleted steps to sharpen the plan’s focus. For example, P12 added a step at the end of their plan to have the agent suggest some potential next steps to pursue after seeing its paper summaries, while P5 removed a less relevant intermediate step on searching for papers related to autonomous vehicles to avoid distractions from this tangential topic. In general, iterative output evaluation and re-planning allowed participants to better “fine-tune” the agent’s outputs and also their own research process—P1 likened their experience to using “[plan steps] as building blocks” that they could mix and match to create a custom workflow.
6.2.2. Fluid movement between co-planning and co-execution enabled granular steering
The interleaving of co-planning and co-execution allowed participants to move fluidly between the two and exert control exactly when necessary. By opting for stepwise execution of the plan, participants could pause the agent to “catch up” with agent activity and ensure outputs are up to standard before continuing. Indeed, P4 and P11 both verified and manually curated the list of papers returned by the agent from a paper search step before proceeding. P11 specifically cites resource efficiency as a practical reason: “If I update [outputs of] the first or second step [after executing all steps], it would need to rerun the following steps. I just want to reduce some API usage.” P10 similarly mentioned that they would “waste energy” running subsequent steps if they later detected errors in initial steps. Overall, many participants chose stepwise execution because it allowed more room to interleave co-planning; this in turn offered more granular control over the human-agent collaborative work trajectory and prevented error propagation.
There were, however, still reasons for preferring continuous execution. P8 and P9 thought continuous execution was more efficient because it parallelized human and AI efforts: “Running all [steps] is a little bit of parallelization in my head where I can get the next thing rolling, and if I end up not having to edit it, that means we’re already moving ahead” (P8). In this case, P8 does not feel a need to be “caught up” with the agent’s work because they know they can still steer the agent at their own pace by modifying the agent’s execution and triggering re-planning at any point.